# Deep Chandra X-ray Observations of Abell 2029: the Merger History of a Relaxed, Strong Cool Core Cluster

### v1

@@ 1 Introduction
One long-standing issue in cluster astrophysics is the cooling flow problem: radiative cooling in the dense cores of relaxed clusters should, in principle, lead to large inflows of gas and high star formation rates in brightest cluster galaxies (BCGs). Yet, observed cooling rates are far lower than predicted (A. C. Fabian, 1994; J. Peterson & A. Fabian, 2006), suggesting some heating mechanism must be involved to balance cooling.
@@
The most likely source of this heating comes from active galactic nuclei (AGN) feedback. The cooling flow of gas that would be deposited in the cluster center, should feed the cluster’s central AGN, resulting in repeated outbursts that can inflate bubbles associated with the AGN’s jets and lobes. These bubbles buoyantly rise to larger radii, distributing heat to the cooling ICM, though the precise mechanism of how this occurs is still debated (B. McNamara & P. Nulsen, 2007; E. L. Blanton et al., 2011). However, this is not the full solution. In some clusters, the central AGN cannot impart enough energy into the surrounding ICM and in these cases, additional processes, such as cluster-cluster interactions which give rise to gas sloshing and shock fronts, can contribute significant heating. While major mergers can permanently disrupt cool cores and raise central entropy beyond the reach of AGN feedback alone, even minor off-axis mergers may produce sloshing motions that help suppress runaway cooling in otherwise relaxed systems (J. A. ZuHone et al., 2010; S.-S. Chen et al., 2024).
@@
Prominent examples of such sloshing include the Perseus Cluster, which hosts a multi-scale spiral structure extending hundreds of kiloparsecs (E. Churazov et al., 2001; S. A. Walker et al., 2017); Abell 2052, which exhibits a sloshing spiral in addition to AGN-inflated bubbles (E. L. Blanton et al., 2003, 2009, 2011); and Abell 119, where deep Chandra imaging reveals a prominent sloshing spiral despite the absence of a traditional cool core (C. B. Watson et al., 2023). In systems such as these, sloshing of the cool, metal-rich ICM core led to contact discontinuities (a.k.a. cold fronts; observed as sharp edges in X-ray brightness) and sometimes merger shocks. At the interfaces of these contact discontinuities, shear flows can lead to the development of Kelvin-Helmholtz Instabilities (KHIs). The growth or suppression of KHI structures depend on the ICM magnetic field strength and viscosity, as shown in various numerical studies (J. A. ZuHone et al., 2010, 2011; J. A. Zuhone et al., 2015; E. Roediger et al., 2012, 2013, 2015).
@@
Abell 2029 (A2029, hereafter) is a (J. Sohn et al., 2019) cool-core cluster, thought to be one of the most relaxed clusters in the universe (D. A. Buote & J. C. Tsai, 1996). Even recently, two studies (XRISM Collaboration, 2025; XRISM Collaboration et al., 2025) using observations from X-ray Imaging and Spectroscopy Mission (XRISM, M. Tashiro et al., 2020) report negligible line-of-sight velocities in the ICM. XRISM Collaboration (2025) find that within the central 180 kpc of the cluster core, the relative velocity of the ICM is at rest with respect to the central BCG, with a 3-sig upper limit of km/s. Building on this, XRISM Collaboration et al. (2025) extend the analysis out to 670 kpc from the cluster core, using three radial pointings covering just the NNW region of the cluster. Their findings also show low bulk motions of the ICM, however they note significant blueshifts in the northern regions suggesting increased bulk motions outside the cluster core. Both XRISM studies agree that the non-thermal pressure fraction is less than 3% of the total pressure, consistent with a relaxed cluster. A third XRISM study (A. Sarkar et al., 2025) using Resolve line-ratio diagnostics reports clear multi-temperature structure in A2029, with cooler (3–4 keV) components superposed on the hot (7–8 keV) ICM in the core and inner northern regions. These findings independently support the presence of cooling phases and azimuthal thermal complexity implicated by the sloshing features we map with Chandra.
@@
Despite this, previous X-ray observations have revealed a large spiral in the ICM consistent with sloshing motions, suggesting recent or on-going merger activity (T. E. Clarke et al., 2004a; R. Paterno-Mahler et al., 2013) occurring in A2029. In addition, M. S. Mirakhor et al. (2022) found that A2029 is possibly connected to a neighboring cluster in the north, Abell 2033 (z=0.0812 (J. Sohn et al., 2019); A2033, hereafter) by a bridge of X-ray gas (significant at ) that is consistent with a Mpc wide filament. Further, J. Sohn et al. (2019) showed that A2033, in addition to a secondary infalling group which they label ’SIG’ (; located south of A2029), are both gravitationally bound to A2029 and would be accreted within the next few Gyr.
@@ 2 Chandra X-ray Observations of A2029
A total of 24 Chandra pointings were used, totaling 515 ksec raw exposure, comprised of a set of two previous ACIS-S observations (ObsIDs 891 & 4977) from Cycle 1, one previous ACIS-I observation (ObsID 6101) from Cycle 3, and 21 new ACIS-I observations (ObsIDs 2XXXX; see Tab. 1) from Cycle 23. The data were processed using CIAO (v4.17; A. Fruscione et al., 2006) and CALDB (v4.12.0) provided by the Chandra X-ray Center (CXC). Level 2 event files were created using chandra_repro, accounting for the given observations’ data mode (FAINT or VFAINT).
@@ 3 Imaging Analysis
The right panel of Fig. 1 shows the diffuse emission in the inner area. A foreground, edge-on spiral galaxy, seen in absorption (T. E. Clarke et al., 2004b), is highlighted with a white arrow and this region is excluded from all detailed analyses presented in this work.
Figure 2 shows an optical RGB image of A2029 obtained from the Legacy Survey Sky Browser333www.legacysurvey.com/viewer, comprised of SDSS g-, r-, and z-band images. Overlaid are the X-ray point sources identified via wavdetect (red ellipses). Also overlaid are cyan contours of 1.4 GHz radio emission, taken from the Very Large Array (VLA) Faint Images of the Radio Sky at Twenty-cm (FIRST) survey, with a beam which has a FWHM size of (R. H. Becker et al., 1995). Contours are shown at levels of 0.0005, 0.0036, 0.0129, 0.0283, and 0.05 Jy beam-1, smoothed with a Gaussian kernel of width 4 pixels. Two additional radio sources south of the central bent-lobe AGN are shown in inset thumbnails (panels (b)-(c) in Fig. 2). We discuss these sources and their relation to the observed ICM substructure in more detail in §5.3 and §5.5.
@@ 3.1 ICM Substructure
To highlight substructure that may not be readily seen in the source image, we fit a 2D -model (beta2d model in Sherpa (v4.16.0; A. Siemiginowska et al., 2024)) to the unsmoothed diffuse emission image in the 0.7-7 keV range using a circular region with radius 123. Corrections for background emission and instrumental exposure were made using the merged blank-sky background image and combined exposure map. The center position, core radius, ellipticity, position angle, amplitude, and power law index of the beta-model were allowed to vary while the amplitudes of the supplied background image and exposure map were frozen. Errors were computed with Cash statistics.
@@
The fitted 2D -model is subtracted from the unsmoothed X-ray image, and the resulting residual image is shown in Fig. 3. The residual emission image was smoothed with a 75 Gaussian, covers an area of 25′25′, and includes arrows pointing to some notable features. The sloshing ICM is easily identified with its spiral-like morphology, spiraling anti-clockwise towards the cluster center (or clockwise outwards). Notably, the spiral excess appears to wrap back to the N, a feature not observed in earlier, shallower Chandra studies (T. E. Clarke et al., 2004a; R. Paterno-Mahler et al., 2013). The spiral excess extends outwards to kpc (excluding the southeastern splash feature), while the SE “splash” (white arrows in Fig. 3) extends kpc beyond the cluster core.
@@ 3.1.2 Gaussian Gradient Magnitude Filter
A Gaussian Gradient Magnitude (GGM) filter was applied to the X-ray images using the gaussian_gradient_magnitude function from SCIPY (P. Virtanen et al., 2020). This filtering method computes the gradients in the two-dimensional data under the assumption of Gaussian derivatives, thereby enhancing the visibility of surface brightness edges (J. S. Sanders et al., 2016a). The key parameter in this technique is , which defines the width of the Gaussian kernel; smaller values emphasize fine-scale structures, whereas larger values reveal more extended features.
@@
Fig. 5 shows a GGM filtered image, zoomed into the inner area, with the 1.4 GHz radio emission contours overlaid in cyan. The WAT lobes show some evidence of being swept back, following the sloshing motions towards the cluster core. This is consistent with what was previously seen in T. E. Clarke et al. (2004a) and R. Paterno-Mahler et al. (2013).
@@ 4 Spectral Analysis
Extracted spectra are fit in the 0.7-7 keV range using XSPEC (v12.14.1) (K. A. Arnaud, 1996). A corrective factor that depends on the observation specific ratio of exposure times and count rates in the hard-band (10-12 keV) range was applied to the observation specific background spectrum before subtracting it from the corresponding source spectrum. We adopted the N. Grevesse & A. J. Sauval (1998) (grsa) solar abundance table, unless otherwise noted. Since spectra are binned to 1 count per bin, all fits are done using C-statistics. However, to report an estimate of the goodness of fit, we take the best-fit models and apply them to spectra that were re-binned to 40 counts/bin and do not re-fit. The resulting is what is reported here. For each spectral fit, the neutral hydrogen column density was fixed at the Galactic value, cm-2 (J. M. Dickey, 1990), and the redshift was fixed to (J. Sohn et al., 2019).
@@ 4.1 Global Spectral Properties
To compare with previous studies, we extract and fit spectra, with weighted responses from the inner 169 kpc, corresponding to the 116″region of C. L. Sarazin & R. W. O’Connell (1993), adjusted for our cosmology. The temperature within this region has been measured, using the A2029 ACIS-S observations, by T. E. Clarke et al. (2004a) as keV and by R. Paterno-Mahler et al. (2013) as keV, where the errors reported by both studies are 90% confidence ranges. Using the same region, we obtain a single-temperature APEC fit of keV and with , consistent with T. E. Clarke et al. (2004a) and A. Vikhlinin et al. (2005) and slightly below R. Paterno-Mahler et al. (2013). Differences in and abundance tables (they used E. Anders & N. Grevesse (1989)) likely contribute. We emphasize that the statistical uncertainties on the projected temperature and abundance are significantly smaller than the systematic uncertainties introduced by our scaling factors (Appendix A). Thus, small differences in the assumed scaling factor could produce larger shifts than the quoted statistical errors.
We have already noted a complex sloshing dynamic in the ICM; thus, there could be more than one temperature component of the gas in the fitted region. Fitting a double APEC model gives best fit parameters of keV, keV, , and . The appropriate scale factors were applied to the upper and lower temperatures and the abundance parameter of the new observations when tied to the fit parameters of the pre-2005 observations, following Appendix A. Running an F-test to compare the single and double APEC fits shows significant improvement with the second APEC component, with an F-statistic of 45.3 and probability . In comparison, with the older, pre-2005 dataset, R. Paterno-Mahler et al. (2013) find best fit values of keV and keV and abundance , where they give the 90% confidence ranges. Thus, we are able to obtain much tighter constraints on the upper and lower confidence of the component.
In the case of A2029, the mass deposition rate, , is still debated. C. L. Sarazin & R. W. O’Connell (1993) and C. B. Peres et al. (1998) used ROSAT imaging to estimate a mass deposition rate within the central 116″ and found that rates of M⊙ yr-1 (scaled to our cosmology) were needed to match observations. From spectra extracted in the same region, R. Paterno-Mahler et al. (2013) used the old Chandra ACIS-S observations (ObsID 891 & 4977) and measured an even higher rate of M⊙ yr-1, cooling over a range of to keV. In contrast, other studies find much lower values — e.g., M⊙ yr-1 (T. E. Clarke et al., 2004a); however, this value is for ICM cooling to very low temperatures, 0.08 keV. Our analysis methods are most comparable to R. Paterno-Mahler et al. (2013) and T. E. Clarke et al. (2004a), who both fit spectroscopic cooling flow models, allowing for a more direct comparison of results.
@@
When allowing both the upper and lower temperatures of the model to vary, we find keV and keV, , and M⊙ yr-1 with . We are consistent with what was seen in R. Paterno-Mahler et al. (2013) where we see a high with cooling over a limited temperature range, down to a lower temperature of 3.7 keV. When is fixed at 0.0808 keV, we find keV, , and M⊙ yr-1 with . This scenario is more consistent with what was seen in T. E. Clarke et al. (2004a) with a much lower of M⊙ yr-1, but cooling all the way down to low temperatures (0.08 keV). The fit with the higher and cutoff in the low temperature of 3.7 keV, however, is more preferred based on the fit statistics.
@@
Previous high–resolution grating work with XMM-Newton/RGS (H. Liu et al., 2019) placed weak constraints on gas cooling below 1 keV in A2029, with upper limits on mass-deposition rates far below the classical cooling-flow prediction. Our Chandra data reveal significant emission from gas at intermediate temperatures of 3–4 keV, implying significant gas cooling down to these levels. Thus our results extend the earlier RGS constraints by demonstrating that while cooling to sub-keV gas remains strongly suppressed, substantial amounts of gas may be cooling down to a few keV within the core. Note that cooling to 1/2 to 1/3 of the starting temperature is seen very often in cooling flows (e.g., A2052; E. L. Blanton et al., 2001, 2003, 2009, 2011).
@@
| (169 kpc) regionaaFollowing C. L. Sarazin & R. W. O’Connell (1993); T. E. Clarke et al. (2004a); R. Paterno-Mahler et al. (2013) | ||||||

### v2

@@ 1 Introduction
One long-standing issue in cluster astrophysics is the cooling flow problem: radiative cooling in the dense cores of relaxed clusters should, in principle, lead to large inflows of gas and high star formation rates in brightest cluster galaxies (BCGs). Yet, observed cooling rates are far lower than predicted (Fabian2003; Peterson2006), suggesting some heating mechanism must be involved to balance cooling.
@@
The most likely source of this heating comes from active galactic nuclei (AGN) feedback. The cooling flow of gas that would be deposited in the cluster center, should feed the cluster’s central AGN, resulting in repeated outbursts that can inflate bubbles associated with the AGN’s jets and lobes. These bubbles buoyantly rise to larger radii, distributing heat to the cooling ICM, though the precise mechanism of how this occurs is still debated (McNamara2007; Blanton2011). However, this is not the full solution. In some clusters, the central AGN cannot impart enough energy into the surrounding ICM and in these cases, additional processes, such as cluster-cluster interactions which give rise to gas sloshing and shock fronts, can contribute significant heating. While major mergers can permanently disrupt cool cores and raise central entropy beyond the reach of AGN feedback alone, even minor off-axis mergers may produce sloshing motions that help suppress runaway cooling in otherwise relaxed systems (ZuHone2010; Chen2024).
@@
Prominent examples of such sloshing include the Perseus Cluster, which hosts a multi-scale spiral structure extending hundreds of kiloparsecs (Churazov2001; Walker2017); Abell 2052, which exhibits a sloshing spiral in addition to AGN-inflated bubbles (Blanton2003; Blanton2009; Blanton2011); and Abell 119, where deep Chandra imaging reveals a prominent sloshing spiral despite the absence of a traditional cool core (Watson2023). In systems such as these, sloshing of the cool, metal-rich ICM core led to contact discontinuities (a.k.a. cold fronts; observed as sharp edges in X-ray brightness) and sometimes merger shocks. At the interfaces of these contact discontinuities, shear flows can lead to the development of Kelvin-Helmholtz Instabilities (KHIs). The growth or suppression of KHI structures depend on the ICM magnetic field strength and viscosity, as shown in various numerical studies (ZuHone2010; ZuHone2011; ZuHone2015; Roediger2012; Roediger2013a; Roediger2015).
@@
Abell 2029 (A2029, hereafter) is a (Sohn2019) cool-core cluster, thought to be one of the most relaxed clusters in the universe (Buote1996). Even recently, two studies (XRISMCollaboration2025; XRISMCollaboration2025a) using observations from X-ray Imaging and Spectroscopy Mission (XRISM, Tashiro2020) report negligible line-of-sight velocities in the ICM. XRISMCollaboration2025 find that within the central 180 kpc of the cluster core, the relative velocity of the ICM is at rest with respect to the central BCG, with a 3-sig upper limit of km/s. Building on this, XRISMCollaboration2025a extend the analysis out to 670 kpc from the cluster core, using three radial pointings covering just the NNW region of the cluster. Their findings also show low bulk motions of the ICM, however they note significant blueshifts in the northern regions suggesting increased bulk motions outside the cluster core. Both XRISM studies agree that the non-thermal pressure fraction is less than 3% of the total pressure, consistent with a relaxed cluster. A third XRISM study (Sarkar2025) using Resolve line-ratio diagnostics reports clear multi-temperature structure in A2029, with cooler (3–4 keV) components superposed on the hot (7–8 keV) ICM in the core and inner northern regions. These findings independently support the presence of cooling phases and azimuthal thermal complexity implicated by the sloshing features we map with Chandra.
@@
Despite this, previous X-ray observations have revealed a large spiral in the ICM consistent with sloshing motions, suggesting recent or on-going merger activity (Clarke2004; PM2013) occurring in A2029. In addition, Mirakhor2022 found that A2029 is possibly connected to a neighboring cluster in the north, Abell 2033 (z=0.0812 (Sohn2019); A2033, hereafter) by a bridge of X-ray gas (significant at ) that is consistent with a Mpc wide filament. Further, Sohn2019 showed that A2033, in addition to a secondary infalling group which they label ’SIG’ (; located south of A2029), are both gravitationally bound to A2029 and would be accreted within the next few Gyr.
@@ 2 Chandra X-ray Observations of A2029
A total of 24 Chandra pointings were used, totaling 515 ksec raw exposure, comprised of a set of two previous ACIS-S observations (ObsIDs 891 & 4977) from Cycle 1, one previous ACIS-I observation (ObsID 6101) from Cycle 3, and 21 new ACIS-I observations (ObsIDs 2XXXX; see Tab. 1) from Cycle 23. The data were processed using CIAO (v4.17; Fruscione2006) and CALDB (v4.12.0) provided by the Chandra X-ray Center (CXC). Level 2 event files were created using chandra_repro, accounting for the given observations’ data mode (FAINT or VFAINT).
@@ 3 Imaging Analysis
The right panel of Fig. 1 shows the diffuse emission in the inner area. A foreground, edge-on spiral galaxy, seen in absorption (Clarke2004b), is highlighted with a white arrow and this region is excluded from all detailed analyses presented in this work.
Figure 2 shows an optical RGB image of A2029 obtained from the Legacy Survey Sky Browser333www.legacysurvey.com/viewer, comprised of SDSS g-, r-, and z-band images. Overlaid are the X-ray point sources identified via wavdetect (red ellipses). Also overlaid are cyan contours of 1.4 GHz radio emission, taken from the Very Large Array (VLA) Faint Images of the Radio Sky at Twenty-cm (FIRST) survey, with a beam which has a FWHM size of (Becker1995). Contours are shown at levels of 0.0005, 0.0036, 0.0129, 0.0283, and 0.05 Jy beam-1, smoothed with a Gaussian kernel of width 4 pixels. Two additional radio sources south of the central bent-lobe AGN are shown in inset thumbnails (panels (b)-(c) in Fig. 2). We discuss these sources and their relation to the observed ICM substructure in more detail in §5.3 and §5.5.
@@ 3.1 ICM Substructure
To highlight substructure that may not be readily seen in the source image, we fit a 2D -model (beta2d model in Sherpa (v4.16.0; Siemiginowska2024)) to the unsmoothed diffuse emission image in the 0.7-7 keV range using a circular region with radius 123. Corrections for background emission and instrumental exposure were made using the merged blank-sky background image and combined exposure map. The center position, core radius, ellipticity, position angle, amplitude, and power law index of the beta-model were allowed to vary while the amplitudes of the supplied background image and exposure map were frozen. Errors were computed with Cash statistics.
@@
The fitted 2D -model is subtracted from the unsmoothed X-ray image, and the resulting residual image is shown in Fig. 3. The residual emission image was smoothed with a 75 Gaussian, covers an area of 25′25′, and includes arrows pointing to some notable features. The sloshing ICM is easily identified with its spiral-like morphology, spiraling anti-clockwise towards the cluster center (or clockwise outwards). Notably, the spiral excess appears to wrap back to the N, a feature not observed in earlier, shallower Chandra studies (Clarke2004; PM2013). The spiral excess extends outwards to kpc (excluding the southeastern splash feature), while the SE “splash” (white arrows in Fig. 3) extends kpc beyond the cluster core.
@@ 3.1.2 Gaussian Gradient Magnitude Filter
A Gaussian Gradient Magnitude (GGM) filter was applied to the X-ray images using the gaussian_gradient_magnitude function from SCIPY (scipy2020). This filtering method computes the gradients in the two-dimensional data under the assumption of Gaussian derivatives, thereby enhancing the visibility of surface brightness edges (Sanders2016). The key parameter in this technique is , which defines the width of the Gaussian kernel; smaller values emphasize fine-scale structures, whereas larger values reveal more extended features.
@@
Fig. 5 shows a GGM filtered image, zoomed into the inner area, with the 1.4 GHz radio emission contours overlaid in cyan. The WAT lobes show some evidence of being swept back, following the sloshing motions towards the cluster core. This is consistent with what was previously seen in Clarke2004 and PM2013.
@@ 4 Spectral Analysis
Extracted spectra are fit in the 0.7-7 keV range using XSPEC (v12.14.1) (Arnaud1996). A corrective factor that depends on the observation specific ratio of exposure times and count rates in the hard-band (10-12 keV) range was applied to the observation specific background spectrum before subtracting it from the corresponding source spectrum. We adopted the Grevesse1998 (grsa) solar abundance table, unless otherwise noted. Since spectra are binned to 1 count per bin, all fits are done using C-statistics. However, to report an estimate of the goodness of fit, we take the best-fit models and apply them to spectra that were re-binned to 40 counts/bin and do not re-fit. The resulting is what is reported here. For each spectral fit, the neutral hydrogen column density was fixed at the Galactic value, cm-2 (Dickey1990), and the redshift was fixed to (Sohn2019).
@@ 4.1 Global Spectral Properties
To compare with previous studies, we extract and fit spectra, with weighted responses from the inner 169 kpc, corresponding to the 116″region of Sarazin1992, adjusted for our cosmology. The temperature within this region has been measured, using the A2029 ACIS-S observations, by Clarke2004 as keV and by PM2013 as keV, where the errors reported by both studies are 90% confidence ranges. Using the same region, we obtain a single-temperature APEC fit of keV and with , consistent with Clarke2004 and Vikhlinin2005 and slightly below PM2013. Differences in and abundance tables (they used Anders1989) likely contribute. We emphasize that the statistical uncertainties on the projected temperature and abundance are significantly smaller than the systematic uncertainties introduced by our scaling factors (Appendix A). Thus, small differences in the assumed scaling factor could produce larger shifts than the quoted statistical errors.
We have already noted a complex sloshing dynamic in the ICM; thus, there could be more than one temperature component of the gas in the fitted region. Fitting a double APEC model gives best fit parameters of keV, keV, , and . The appropriate scale factors were applied to the upper and lower temperatures and the abundance parameter of the new observations when tied to the fit parameters of the pre-2005 observations, following Appendix A. Running an F-test to compare the single and double APEC fits shows significant improvement with the second APEC component, with an F-statistic of 45.3 and probability . In comparison, with the older, pre-2005 dataset, PM2013 find best fit values of keV and keV and abundance , where they give the 90% confidence ranges. Thus, we are able to obtain much tighter constraints on the upper and lower confidence of the component.
In the case of A2029, the mass deposition rate, , is still debated. Sarazin1992 and Peres1998 used ROSAT imaging to estimate a mass deposition rate within the central 116″ and found that rates of M⊙ yr-1 (scaled to our cosmology) were needed to match observations. From spectra extracted in the same region, PM2013 used the old Chandra ACIS-S observations (ObsID 891 & 4977) and measured an even higher rate of M⊙ yr-1, cooling over a range of to keV. In contrast, other studies find much lower values — e.g., M⊙ yr-1 (Clarke2004); however, this value is for ICM cooling to very low temperatures, 0.08 keV. Our analysis methods are most comparable to PM2013 and Clarke2004, who both fit spectroscopic cooling flow models, allowing for a more direct comparison of results.
@@
When allowing both the upper and lower temperatures of the model to vary, we find keV and keV, , and M⊙ yr-1 with . We are consistent with what was seen in PM2013 where we see a high with cooling over a limited temperature range, down to a lower temperature of 3.7 keV. When is fixed at 0.0808 keV, we find keV, , and M⊙ yr-1 with . This scenario is more consistent with what was seen in Clarke2004 with a much lower of M⊙ yr-1, but cooling all the way down to low temperatures (0.08 keV). The fit with the higher and cutoff in the low temperature of 3.7 keV, however, is more preferred based on the fit statistics.
@@
Previous high–resolution grating work with XMM-Newton/RGS (Liu2019) placed constraints on gas cooling below 1 keV in A2029, with upper limits on mass-deposition rates far below the classical cooling-flow prediction. Our Chandra data reveal significant emission from gas at intermediate temperatures of 3–4 keV, implying significant gas cooling down to these levels. Thus our results extend the earlier RGS constraints by demonstrating that while cooling to sub-keV gas remains strongly suppressed, substantial amounts of gas may be cooling down to a few keV within the core. Note that cooling to 1/2 to 1/3 of the starting temperature is seen very often in cooling flows (e.g., A2052; Blanton2001; Blanton2003; Blanton2009; Blanton2011).
@@
| (169 kpc) regionaaFollowing Sarazin1992; Clarke2004; PM2013 | ||||||